OpenStudy (anonymous):

What is the solution to the equation? 32- 6x + 7x = -15 + 28 x = __

OpenStudy (anonymous):

use mathway.com it's amazing.

OpenStudy (welshfella):

what does -6x + 7x come to ?

OpenStudy (anonymous):

1x

OpenStudy (anonymous):

Does x equal 19?

OpenStudy (welshfella):

ok 1x or x is correct

OpenStudy (welshfella):

32 + x = -15 + 28

OpenStudy (welshfella):

x = -15 + 28 - 32

OpenStudy (anonymous):

if a number and a letter are togather it means for example 2xh

OpenStudy (welshfella):

@WeepingWillow - can you work out the right side of the equation 19 is 'nearly' right

OpenStudy (anonymous):

-19?